N(t)=-3t^3+23t^2+8t
a)Rewrite the formula by factoring the right-hand side completely.
b)Use the factored version of the formula to find N(3).
c)Use the accompanying graph to estimate the time at which the workers are most efficient.
d)Use the accompanying graph to estimate the maximum number of components assembled per hour during an 8-hour shift.

Answer by stanbon(75887)

N(t)=-3t^3+23t^2+8t
a)Rewrite the formula by factoring the right-hand side completely.
N(t) = -t(3t^2 -23t -8)
N(t) = -t(3t^2-24t + t-8)
N(t) = -t(3t(t-8) + (t-8))
N(t) = -t(t-8)(3t+1)
============================================================
b)Use the factored version of the formula to find N(3)
N(3) = -3(-5)(10) = 150
============================================================
c)Use the accompanying graph to estimate the time at
which the workers are most efficient.
Graph not shown
=========================
d)Use the accompanying graph to estimate the maximum number of components assembled per hour during an 8-hour shift.
Graph not shown
=========================
Cheers,
Stan H.